A sample of phosphorus has 0.5 moles of P4 molecules, what is the mass of the sample?
1
Expert's answer
2019-08-05T02:08:43-0400

First, the molar weight of the molecular phosphorus P4 is:

"M(P_{4})=4*31g\/mol=124g\/mol"


And now, we can calculate the mass of the phosphorus P4:

"m(P_{4})=n(P_{4})*M(P_{4})=0.5mol*124g\/mol=62g"

Thus, the mass of 0.5 mol of P4 molecules is 62 grams.